OpenStudy (anonymous):

A movie rental store sells a couples movie pack for $6 and a family movie pack for $10 . In one year, the Nguyen family paid $126 for 13 movie packs. How many family movie packs did they pay for?

OpenStudy (anonymous):

Didn't it just say how many movie packs they paid for? ( confused ) :S

OpenStudy (radar):

Let x = number of family packs. Then, 13-x = number of couples movies. The cost would then be: 10x + 6(13-x)=126 (dollars) 10x +78-6x=126 4x=48 x=12 family packs 13-12=1 couples movies

OpenStudy (anonymous):

thanks

OpenStudy (radar):

ur welcome